OpenStudy (anonymous):

There are (2^4)^2 multiply 20 books in the cupboard. What is the total number of books in the cupboard?

OpenStudy (misty1212):

add, use a calculator

OpenStudy (misty1212):

when you add we can write in scientific notation

OpenStudy (misty1212):

whoa the question changed or am i crazy?

OpenStudy (mathstudent55):

You mean this? \((2^4)^2 \times 20\)

OpenStudy (anonymous):

correct

OpenStudy (mathstudent55):

Start working from inside the parentheses. What is \(2^4\)?

OpenStudy (anonymous):

2x2x2x2 right?

OpenStudy (anonymous):

16

OpenStudy (mathstudent55):

Yes. Now the next step. What is \(16^2\) ?

OpenStudy (anonymous):

16x16 = 256

OpenStudy (mathstudent55):

Great. Now you can finish with: \(256 \times 20\)

OpenStudy (anonymous):

256 x 20 = 5120

OpenStudy (anonymous):

I think were doing this wrong

OpenStudy (mathstudent55):

Correct. The way a problem like this should be done is by doing it one step at a time like we did, and copying everything down each step. This way you don't forget any of the steps or operations. See below.

OpenStudy (anonymous):

what about the exponents I cant just multiply smoothly like that because it gives me an answer that has nothing to do with my options

OpenStudy (mathstudent55):

\((2^4)^2 \times 20\) \(= 16^2 \times 20\) \(= 256 \times 20\) \(= 5120\)

OpenStudy (anonymous):

the way I should of done it is to keep my base and add the exponents

OpenStudy (mathstudent55):

We solved together the problem you posted. Perhaps you did not copy the problem correctly. Can you post a pic of the problem or use the draw tool to hand write the exact problem?

OpenStudy (anonymous):

I copy and paste the whole entire question maybe I should do the same for my options no offence though. :)

OpenStudy (mathstudent55):

You add exponents when you multiply two powers with the same base.

OpenStudy (mathstudent55):

I did ask you above if I was interpreting the problem correctly, and you said yes.

OpenStudy (anonymous):

ill copy the question again

OpenStudy (mathstudent55):

Some symbols cannot be simply cut and pasted.

OpenStudy (anonymous):

I solved the problem myself thank you for the help though :)